# An alternative theory of magnetic flux tubes in strong fields via axion origin photons

### v2

@@ I Introduction
Figure 6: A sketch of the magnetic tube born anchored to the tachocline and risen to the solar surface by the neutral buoyancy ($\rho_{ext}=\rho_{int}^{0}$). The strong convection suppression inside the tube leads to the abrupt decrease of temperature and density ($\rho_{int}^{0}\gg\rho_{int}$), which in its turn leads to the significant decrease in gas pressure above the umbra (sunspot). At the top of the overshoot tachocline the first term $(dQ/dt)_{1}$ characterizes the radiative heating which depends on the thermodynamic quantities $k_{e}xt$ and $T_{ext}$ of the external plasma, changing with the distance from the center of the Sun only (see Eq. (20)). The second term $(dQ/dt)_{2}$ represents the diffuse radiation through the flux tube because of the temperature difference between the tube ring (see $a\equiv a_{conv}$ between the O-loop and the tube walls) and the surrounding plasma (see Eq. (21)). The keV photons (see Fig. 2 in J. E. Bailey et al. (2009)) coming from the radiation zone are turned into axions in the horizontal magnetic field of the O-loop (see Fig. 3). Therefore, the radiative heating almost vanishes in the virtually empty magnetic tube. The base of the convection zone is defined as a radius at which the stratification switches from almost adiabatic ($\delta=\nabla_{e}-\nabla_{ad}=0$) to sub-adiabatic ($\delta=\nabla_{e}-\nabla_{ad}<0$). Meanwhile, the external plasma turns from sub-adiabatic to super-adiabatic ($\delta=\nabla_{e}-\nabla_{ad}>0$).
At the same time, we remember that anticorrelated 11-year ADM density modulations in the Sun interior (see Fig. 1b) and the ADM around BH (see Sect. 3 in (V. D. Rusov et al., 2021)) are not only interconnected, but simultaneously, as consequence, are driving the 11-year modulations of the solar axion density (see Eqs. (27)-(28) in (V. D. Rusov et al., 2021)).
As a result, assuming 11-year variations in convective heating $(dQ/dt)_{2}$ in the form of a maximum and minimum of the “ring” width of the $a\equiv a_{conv}$ (between the O-loop and the tube walls in Fig. 6)
$(a_{conv})_{max}\sim 3.7\cdot 10^{-4}H_{p}\approx 22~km,$ (22)
$(a_{conv})_{min}\sim 3.7\cdot 10^{-3}H_{p}\approx 220~km,$ (23)
we apply a new analysis of the universal vanBFF model (see Eqs. (32.1)-(18) analogous to (A. A. van Ballegooijen, 1982) or (19)-(21) (analogous to the last part of Eq. (13) and (19) in (Y. Fan & G. Fisher, 1996))), where the calculated values such as the magnetic flux $\Phi$ and the rise speed $(v_{rise})_{conv}$ of the MFT to the surface of the Sun, do not contradict the known observational data:
$\bullet$ values of the magnetic tube (see Fig. 5a), based on the physics of convective heating $(dQ/dt)_{2}$, are related to the radius of the “ring” between the O-shaped contour and the walls of the magnetic tube (Fig. 6), and which are also the result of maximum magnetic flux
@@
where $r\sim 220~km$, $(a_{conv})_{min}\sim 22~km$ and $(B_{tacho})_{max}\sim 3.6\cdot 10^{7}~G$, and minimum magnetic flux
$\Phi_{min}=\pi(a_{conv})_{max}^{2}(B_{tacho})_{min}\approx 5\cdot 10^{21}~Mx,$ (25)
where $\pi(a_{conv})_{max}^{2}\approx\pi(220^{2}-20^{2})~km^{2}=1.5\cdot 10^{15}~cm^{2}$ and $(B_{tacho})_{min}\sim 3.6\cdot 10^{6}~G$, and which is in good agreement with the observational data of C. Zwaan (1987) (see also Section 2.1 in (Y. Fan, 2021)) since the total magnetic flux ranges are $1-5\cdot 10^{22}~Mx$ (near the maximum) and $2-5\cdot 10^{21}~Mx$ (near the minimum) (see (K. L. Harvey et al., 1982)).
@@
It is very important here that, on the one hand, the anchored magnetic field in the tachocline rises from the bottom of the convective zone to the solar photosphere in the form of thin isolated filaments known as flux tubes (see e.g. (E. N. Parker, 1979a)), and on the other hand, the well-known nonlinear dynamics of a thin magnetic tube was modeled in the form that has been used by H. C. Spruit (1981). In this case, the magnetic tube is assumed to be thin in the sense that its cross-sectional radius $a_{conv}$ is negligibly small with respect to both the atmospheric scale height (i.e. $a_{conv}\ll H_{p}$ (see Fig. 6)) and any variation scales along the tube. This means that each magnetic field line is roughly parallel to the axis of the tube, and the Spruit’s model is applicable only to untwisted magnetic tubes.
@@
This raises a rather complicated but simple question: In what way, without twisting, can a real magnetic tube, which is not integral and will not behave as a single object for a very long time (E. N. Parker, 1979a; H. C. Spruit, 1981; D. W. Longcope & I. Klapper, 1997), be obtained without destruction by hydrodynamic forces?
@@
The answer, oddly enough, is very simple! For an untwisted thin magnetic tube, the pressure balance is always maintained not only along the entire radius of the “ring” between the O-shaped contour and the walls of the magnetic tube (Fig. 6), which is solved without destruction by hydrodynamic forces, but at the same time leads to the suppression of convection along the inner radius of the O-shaped contour due to a sharp decrease in temperature and density (see $\rho_{int}^{0}\gg\rho_{int}$ in Fig. 6). The latter means that in very strong magnetic fields (e.g. $5\cdot 10^{5}~G<B<10^{8}~G$) an untwisted thin magnetic tube will, surprisingly, always be perpendicular to the azimuthal direction from the tachocline to the surface of the photosphere!
@@
The physics of this process is briefly as follows. The high-energy photons from the radiation zone through axion-photon oscillations in the O-loop inside the magnetic tube near the tachocline create the so-called axions of photon origin under the sunspot. This means that at such strong magnetic fields, the Parker-Biermann effect of cooling inside the magnetic tube, where the Sun’s gravitational field pulls gas out of the inner tube (as described by hydrostatic pressure, known as the barometric law $dp/dz=-\rho g$), develops due to the “disappearance” of the convective heat transfer $(dQ/dt)_{2}$ and, as a consequence, the temperature in the lower part of the magnetic tube with the help of axions of photon origin from photon-axion oscillations in the O-loop near the tachocline. As a result, it opens a free path (in the radial direction!) for photons of axion origin (Rosseland length; see Fig. B.3 in (V. D. Rusov et al., 2021)) from the tachocline to the photosphere (see Fig. 5a)!
@@
$\bullet\bullet$ the time of rise of convective heating of the flux tube from the tachocline to the photosphere (see Eq. (15)), which corresponds to the equations
@@
at the maximum value of the 11-year variations of the magnetic flux tube in the photosphere for sunspots (see (27)), in the form
@@
where $p_{ext}=6.53\cdot 10^{13}~erg/cm^{3}$ (see A.16 in (V. D. Rusov et al., 2021), and also (J. Bahcall & M. Pinsonneault, 1992)), $T_{ext}=2.3\cdot 10^{6}~K$ and $\rho_{ext}=0.2~g/cm^{3}$ (J. Bahcall & M. Pinsonneault, 1992); $c_{p}=2.5\Re=2.5p_{ext}/\rho_{ext}T_{ext}$ ($\Re$ is the gas constant in the equation of state $p_{ext}=\rho_{ext}\Re T_{ext}=nk_{B}T_{ext}$ (E. Böhm-Vitense, 1958)); $\mu=0.58$ is mean molecular weight of hydrogen in overshoot tachocline, $m_{p}\cong 1.67\cdot 10^{-24}~g$ is the proton mass: ”2018 CODATA Value: proton mass”, The NIST Reference on Constants, Units, and Uncertainty, 20 May 2019; $k_{B}=1.38\cdot 10^{-23}~J/K$; $(dQ/dt)_{1}\approx 29.7~erg\cdot cm^{-3}\cdot s^{-1}$ (see Eq. (18) in (Y. Fan & G. Fisher, 1996)) is the rate of radiative heating; $\alpha_{1}^{2}\approx 5.76$ (Y. Fan & G. Fisher, 1996; M. A. Weber & Y. Fan, 2015); $\nabla_{e}\approx 0.4$ (see Table 2 in (H. Spruit, 1974)); $H_{p}=\Re T_{ext}/g=p_{ext}/g\rho_{ext}=0.08R_{Sun}$ (E. Böhm-Vitense, 1958; H. C. Spruit, 1977; A. S. Brun et al., 2011); $a_{conv}^{max}\sim 3.7\cdot 10^{-4}H_{p}$ (see Eq. (31)).
In order to obtain the minimum values of the 11-year variations of the magnetic flux tube in the photosphere for sunspots, we have to calculate a new value for $p_{ext}=(nk_{B}T)_{min}$, which together with $a_{conv}\sim 3.7\cdot 10^{-3}~H_{p}$ (Eq. (23)) is related to Eqs. (32)-(33). Using the value of fully ionized hydrogen plasma with $Z=1$ in the tachocline, which is related to the equation $T^{1/4}n=const$ (see Eq. (A.9) from Appendix A in (V. D. Rusov et al., 2021)), we obtain the relation
$(T^{1/4}\rho_{ext})_{min}=(T^{1/4}\rho_{ext})_{max}=((2.3\cdot 10^{6}~K)^{1/4}\times 0.2~g/cm^{3})_{max}.$ (32)
When we apply the temperature “minimum” value $(T\sim 10^{6}~K)_{min}$ (see Eq. (28) from (V. D. Rusov et al., 2021)), then we get the following relation
$\mu\times m_{p}\times(T^{1/4}n)_{min}=((10^{6}~K)^{1/4}\rho_{ext})_{min}=\mu\times m_{p}\times(T^{1/4}n)=((2.3\cdot 10^{6}~K)^{1/4}\times 0.2~g/cm^{3})_{max},$ (33)
from which we obtain the complete solution of the “minimum” rise time of the convective heating of the flux tube from the tachocline to the photosphere (see Eq. (15), which corresponds to the equation
$(\tau_{d})_{conv}^{min}\approx 5.7\cdot 10^{5}~s\sim 7~days,$ (34)
where $p_{ext}=(nk_{B}T)_{min}=3.56\cdot 10^{13}~erg/cm^{3}$, $T_{ext}=10^{6}~K$, $\rho_{ext}\approx 0.25~g/cm^{3}$ (see Eq. 33).
@@
which are almost identical to the observational data of the known works (see about 1.4 m/s at 62-75 and 42-55 Mm for AR11726, April 2013 in A. G. Kosovichev et al. (2016, 2018); see about 1.2 km/s at 16-32 Mm for AR NOAA 6432, South Pole in 1991 January 4-5, in A. Kosovichev (1996)), and at the minimum of the solar cycle
$(v_{rise})_{conv}^{min}\sim 0.25~km/s,$ (37)
which are also almost identical to the observational data, e.g. of the known works by S. Ilonidis et al. (2011, 2012, 2013). This is due to the fact that these works, on the one hand, showed the measured velocities of the magnetic flux escape from the deep interior of the Sun into the atmosphere in the form of sunspots (see $\sim 1~km/s$ for 2003 October 25 (Fig. 5.14 in S. Ilonidis et al. (2012); Fig. 8 in S. Ilonidis et al. (2013))), and on the other hand, using deep-focus time-distance helioseismology T. L. Duvall et al. (1993, 1996); A. G. Kosovichev et al. (2016) found several strong emerging flux events (for AR 10488, October 26, 2003, + ARs 8164 and 8171, 23-27 February 1998 + AR 7978, 06 July 1996) at depths of 42-75 Mm up to 6 h (using only previous data) before entering the photosphere and 1-2 days before until the appearance of the detected magnetic structures on the surface with an average speed of 0.3 to 0.6 km/s (S. Ilonidis et al., 2011). This means that, unlike the works of S. Ilonidis et al. (2011, 2012, 2013), we use the velocities of emerging flows on the surface of sunspots between the maximum ((31) and (36)) and minimum ((34) and (37)) of the solar cycle (for AR 10488, October 26, 2003, + ARs 8164 and 8171, 23-27 February 1998 + AR 7978, 06 July 1996), which correspond close to the maximum $\sim 1~km/s$ (for 2003 October), lower (by assumption!) than middle of the maximum and minimum 0.3 km/s +0.3 km/s (for ARs 8164 and 8171, 23-27 February 1998) and low $\sim 0.25~km/s$ (for AR 7978, 06 July 1996). From here, we get the average value of the rate of rise of the magnetic tube from the maximum (for 2003 October) to the minimum (for July 1996)
$\frac{1.0(2003)+0.30(1998)+0.30(1998)+0.25(1996)}{4}km/s\sim 0.45~km/s$ (38)
which due to our mean velocity uncertainty of $45_{-0.15}^{+0.15}~km/s$ almost exactly corresponds to the known uncertainty in the average velocity from 0.3 to 0.6 km/s (S. Ilonidis et al., 2011)!
In other words, this means that our (maximum (31) and (36)) and minimum ((34) and (37)) are the remarkable existence of a maximum and minimum of 11-year variations of the photosphere magnetic flux tube for sunspots, which, unlike the solar dynamo, are predetermined as flux tube magnetic field sources (anchored in the overshoot tachocline), and are coupled by the holographic mechanism of quantum gravity (see Fig. 1b), and anticorrelated 11-year ADM density variations gravitationally trapped in the interior of the Sun (see Section 3 in (V. D. Rusov et al., 2021)).
In order to show signatures of appearing areas of sunspots inside the Sun before they appear on the surface, we used, according to A. G. Kosovichev et al. (2016), the measurements of plasma flows in the upper convection zone, provided by the Time-Distance Helioseismology pipeline developed for the analysis of solar oscillation data obtained by the Helioseismic and Magnetic Imager (HMI) aboard the Solar Dynamics Observatory (SDO), to investigate the subsurface dynamics of the emerging active region (AR) NOAA 11726.
Figure 7c shows the distribution of the effective phase shift (“helioseismic index”) for the depth range 62-75 Mm measured on April 19, 2013, 03:00 UT, when there was no significant magnetic flux on the surface. The figure 7d shows the variations of the helioseismic perturbation associated with the emerging AR at two different depths: 62-75 Mm and 42-55 Mm depending on time. One of the greatest advantages of the remote helioseismology is that the full-disk observations of solar variations allow to build a special scheme of measurement and selection of acoustic waves signals (Fig. 7b), passing though certain sub-surface areas. For example, S. Ilonidis et al. (2013) developed a special procedure of deep focusing which is able to detect large emerging active areas more than a day before their appearance on the surface (Fig. 7). The rise velocity determined by tracking the magnetic flux declination with depth, is about 1.4 km/s, which is very close to the rise velocity in the deep layers A. G. Kosovichev et al. (2016), e.g. near tachocline! Interestingly, the results of several attempts to detect the emerging magnetic flux before its appearance in the photosphere may be compared to our theoretical estimates (maximum (31) and (36)) and minimum ((34) and (37)) and demonstrate a surprising agreement in terms of the magnetic tube rising velocity and time.
Figure 7: Detection of emerging active region AR 11726: a) magnetograms of helioseismic and magnetic imager (HMI) on board of the Solar Dynamics Observatory (SDO) of April 19, 2013, 3:00 UT, prior the emergence; the rings show the area where the acoustic oscillation signal was measured to detect the subsurface signal at the central point; b) a vertical cut through the computational grid used in the time-distance inversions, and a sample of ray paths of acoustic waves, which were used for measuring travel times (A. G. Kosovichev & J. Duvall, 2008; T. L. Duvall et al., 1993, 1996); c) variations of a travel-time index showing the AR perturbation located the depth of 62-75 Mm on April 19, 2013, 3:00 UT; d) variations of the helioseismic perturbation associated with the emerging AR at two different depths: 42-55 Mm and 62-75 Mm, as a function of time. The characteristic rise time estimated using the delays in two layers at depths 42-55 Mm and 62-75 Mm is about 1.4 km/s. Derived from A. G. Kosovichev et al. (2016).
Below we briefly and simply show how our fundamental physics of the rise of strong magnetic flux tubes with $B\gg 10^{5}~G$, but without dynamo action, differs from the classical physics of an active region, in which the rise of magnetic tubes from a tachocline to the photosphere is determined by the result of the convective action of the dynamo in the main part of the solar convection zone (with 0.15-0.25 km/s; see e.g. Y. Fan (2021); P. Charbonneau (2020); P. Charbonneau & D. Sokoloff (2023) and Fig. 21 of H. Hotta & H. Iijima (2020).
Let us first draw some attention to what the remarkable physicist Yuhong Fan says about the essence of magnetic fields in the solar convection zone: “The question of whether, or to what extent, a strong toroidal magnetic field stored in the overshoot region at the base of the convection zone, generated by a deep seated solar dynamo process, is responsible for the formation of solar active regions remains to be more rigorously investigated.(Y. Fan, 2021)
Recalling the words of Yuhong Fan that “it remains to be more thoroughly investigated” to what extent is strong toroidal magnetic field, control by means of the self-generating dynamo in the overshoot tachocline region, one must also remember the existence of the amazing 11-year variation of dark matter density (see e.g. V. D. Rusov et al. (2021)), and, for example, the holographic mechanism of quantum gravity (see S. W. Hawking (2015a, b); S. W. Hawking et al. (2016); V. D. Rusov et al. (2021)), which strongly influence the Sun, the stars, the Universe?! This means that, unlike the holographic mechanism of quantum gravity on the Sun, where the magnetic flux tubes in strong fields ($\sim 4.1\cdot 10^{7}~G$ in the tachocline are identical to the magnetic core of the Sun), such physics of formation, emergence and “decay” of sunspots, which is considered an integral part of the convective dynamo (see e.g. Y. Fan & F. Fang (2014) and Fig. 21 in H. Hotta & H. Iijima (2020)), and, as a consequence, the formation of active regions from the uppermost convection zone to the corona (see e.g. F. Chen et al. (2017, 2022)), is still an unsolved problem of the 70-year evolution and the current state of solar and stellar dynamo theory (see e.g. P. Charbonneau & D. Sokoloff (2023) and P. Charbonneau (2020)), which, although good, cannot be on the Sun, and certainly not in our Universe!
In contrast to the classical physics of rise of magnetic flux tubes from the tachocline to the photosphere, which are driven by a self-generating dynamo in the main part of the solar convective zone, we already know that,
  • first, a strong toroidal magnetic field stored in the overshoot tachocline region is determined by the thermomagnetic Ettingshausen-Nernst effect (see Apendix A in V. D. Rusov et al. (2021)), which exactly (in contrast to dynamo!) coincides with $B_{tacho}\sim 10^{7}~G$. In this case, we indirectly showed that, using the fundamental holographic principle of quantum gravity (see Apendix C in V. D. Rusov et al. (2021)), the repelling toroidal magnetic field of the tachocline exactly “neutralizes” the magnetic field in the Sun’s core (see Eq. 1 and W. A. Fowler et al. (1955); S. Couvidat et al. (2003)), in which the projections of the magnetic fields of the tachocline and the core have the equal value but the opposite directions.
  • second, we underline that this process is associated with continuous transformation of toroidal magnetic field into poloidal magnetic field ($T\rightarrow P$ transformation), but not vice versa ($P\rightarrow T$). It means that the holographic mechanism is the main process of regeneration of the primary toroidal field in the tachocline, and the formation of the processes of these fields shears at the base of the convective zone leads to the emergence of buoyancy of toroidal magnetic tubes rising to the surface of the Sun.
  • third, our physics of the magnetic flux tubes rise is related to the fact that if the axions, which are born in the Sun’s core, are directly converted into X-rays near the tachocline, then the axion-photon oscillations predetermine the appearance of magnetic sunspot cycles. This is due to the fact that the formation of sunspots and their cycles is a consequence, according to V. D. Rusov et al. (2021), of the anticorrelation 11-year cycles of ADM density modulation inside the Sun, which, not surprisingly, are a consequence of the 11-year ADM halo density modulation in the fundamental plane of the galactic center, which closely correlates with the density modulation of baryonic matter near the supermassive black hole. From here, it is easy to show (see V. D. Rusov et al. (2021)) in what way the anticorrelation identity between indicators of the ADM density modulation inside the Sun and the number of sunspots (or the correlation identity between indicators of modulation of solar axions (or photons of axion origin) and sunspot cycles) is realized.
Finally, let us show that if the rise velocity of the buoyant magnetic tube, determined by the magnetic field values maximum ($B_{max}\sim 3.6\cdot 10^{7}~G$) and minimum ($B_{max}\sim 3.6\cdot 10^{6}~G$) of the magnetic field and the width of the “ring” $(a_{conv})_{max}\sim 3.7\cdot 10^{-4}~H_{p}$ or $(a_{conv})_{min}\sim 3.7\cdot 10^{-3}~H_{p}$, causes the MFT to appear on the surface in the form of sunspots, then the parameters of the universal vanBFF model will be associated with the magnetic cycles that are directly related to the observed features of the cycle of both double sunspot maxima (see Sect. II.2.2) and the tilt angle of Joy’s law (see Sect. II.2.1).

### v3

@@ II Physics of the theory of practically empty flux tubes in strong fields of the order $\sim 10^{7}~G$
Figure 5: A sketch of the magnetic tube born anchored to the tachocline and risen to the solar surface by the neutral buoyancy ($\rho_{ext}=\rho_{int}^{0}$). The strong convection suppression inside the tube leads to the abrupt decrease of temperature and density ($\rho_{int}^{0}\gg\rho_{int}$), which in its turn leads to the significant decrease in gas pressure above the umbra (sunspot). At the top of the overshoot tachocline the first term $(dQ/dt)_{1}$ characterizes the radiative heating which depends on the thermodynamic quantities $k_{ext}$ and $T_{ext}$ of the external plasma, changing with the distance from the center of the Sun only (see Eq. (20)). The second term $(dQ/dt)_{2}$ represents the diffuse radiation through the flux tube because of the temperature difference between the tube ring (see $a\equiv a_{conv}$ between the O-loop and the tube walls) and the surrounding plasma (see Eq. (21)). The keV photons (see Fig. 2 in [73]) coming from the radiation zone are turned into axions in the horizontal magnetic field of the O-loop (see Fig. 2). Therefore, the radiative heating almost vanishes in the virtually empty magnetic tube. The base of the convection zone is defined as a radius at which the stratification switches from almost adiabatic ($\delta=\nabla_{e}-\nabla_{ad}=0$) to sub-adiabatic ($\delta=\nabla_{e}-\nabla_{ad}<0$). Meanwhile, the external plasma turns from sub-adiabatic to super-adiabatic ($\delta=\nabla_{e}-\nabla_{ad}>0$).
At the same time, we remember that anticorrelated 11-year ADM density modulations in the Sun interior and the ADM around BH (see Sect. 3 in [14]) are not only interconnected, but simultaneously, as consequence, are driving the 11-year modulations of the solar axion density (see Eqs. (27)-(28) in [14]).
As a result, assuming 11-year variations in convective heating $(dQ/dt)_{2}$ in the form of a maximum and minimum in the width of the “ring” $a\equiv a_{conv}$ (between the O-loop and the tube walls in Fig. 5), we apply a new analysis of the universal vanBFF model (see Eqs. (19)-(21) similar to the last part of Eqs. (13) and (19) in [64])), where the values such as the magnetic flux $\Phi$ (see [71, 72]) and the magnetic fields at the tachocline ($3.6\cdot 10^{7}~G$ in Fig. 2) are consistent with known observational data:
$\bullet$ values of the magnetic tube (see Fig. 4a), based on the physics of convective heating $(dQ/dt)_{2}$, are related to the radius of the “ring” between the O-shaped loop and the walls of the magnetic tube (Fig. 5), and correspond exactly to the “observed” values of the ratio between the maximum and minimum magnetic flux and the magnetic field at the tachocline [71, 72, 69]. Therefore, we are interested in the observed magnetic fluxes
@@
where $\Phi_{max}\approx 10^{22}~Mx$ is at the maximum of the “measured” magnetic fluxes [71, 72] and the maximum of the “measured” magnetic field MFT $(B_{tacho})_{max}\sim 3.7\cdot 10^{7}~G$ (see Fig. 2), which is associated with the so-called Ettingshausen-Nernst thermomagnetic effect, caused by the Parker-Biermann cooling effect.
Then a minimum of magnetic flux appears
$\Phi_{min}=2\pi r(a_{conv})_{max}(B_{tacho})_{min}\approx 5\cdot 10^{21}~Mx,$ (23)
@@
where $\Phi_{min}\approx 5\cdot 10^{21}~Mx$ is at the minimum of the “measured” magnetic fluxes [71, 72] and the minimum of the “measured” magnetic field MFT $(B_{tacho})_{min}\sim 3.7\cdot 10^{6}~G$.
@@
It is very important here that, on the one hand, the anchored magnetic field in the tachocline rises from the bottom of the convective zone to the solar photosphere in the form of thin isolated filaments known as flux tubes (see e.g. [56]), and on the other hand, the well-known nonlinear dynamics of a thin magnetic tube was modeled in the form that has been used by [59]. In this case, the magnetic tube is assumed to be thin in the sense that its cross-sectional radius $a_{conv}$ is negligibly small with respect to both the atmospheric scale height (i.e. $a_{conv}\ll H_{p}$ (see Fig. 5)) and any variation scales along the tube. This means that each magnetic field line is roughly parallel to the axis of the tube, and the Spruit’s model is applicable only to untwisted magnetic tubes.
@@
This raises a rather complicated but simple question: In what way, without twisting, can a real magnetic tube, which is not integral and will not behave as a single object for a very long time [56, 59], be obtained without destruction by hydrodynamic forces?
@@
The answer, oddly enough, is very simple! For an untwisted thin magnetic tube, the pressure balance is always maintained not only along the entire radius of the “ring” between the O-shaped contour and the walls of the magnetic tube (Fig. 5), which is solved without destruction by hydrodynamic forces, but at the same time leads to the suppression of convection along the inner radius of the O-shaped contour due to a sharp decrease in temperature and density (see $\rho_{int}^{0}\gg\rho_{int}$ in Fig. 5). The latter means that in very strong magnetic fields (e.g. $5\cdot 10^{5}~G<B<10^{7}~G$) an untwisted thin magnetic tube will, surprisingly, always be perpendicular to the azimuthal direction from the tachocline to the surface of the photosphere!
@@
The physics of this process is briefly as follows. The high-energy photons from the radiation zone through axion-photon oscillations in the O-loop inside the magnetic tube near the tachocline create the so-called axions of photon origin under the sunspot. This means that at such strong magnetic fields, the Parker-Biermann effect of cooling inside the magnetic tube, where the Sun’s gravitational field pulls gas out of the inner tube (as described by hydrostatic pressure, known as the barometric law $dp/dz=-\rho g$), develops due to the “disappearance” of the convective heat transfer $(dQ/dt)_{2}$ and, as a consequence, the temperature in the lower part of the magnetic tube with the help of axions of photon origin from photon-axion oscillations in the O-loop near the tachocline. As a result, it opens a free path (in the radial direction!) for photons of axion origin (Rosseland length; see Fig. B.3 in [14]) from the tachocline to the photosphere (see Fig. 4a)!
From here, using Eq. (22) and Eq. (23), we obtain the cross-section values $a_{conv}$ (between the O-loop and the tube walls in Fig. 4a) at the minimum
$(a_{conv})_{min}\sim 22~km\approx 3.7\cdot 10^{-4}H_{p}$ (24)
and at the maximum
$(a_{conv})_{max}\sim 220~km\approx 3.7\cdot 10^{-3}H_{p},$ (25)
@@
at which the rise time of convective heating of the flux tube from the tachocline to the photosphere (see Eq. (15)) corresponds to the equations
@@
where at the maximum the values of $(a_{conv})_{min}\approx 3.7\cdot 10^{-4}H_{p}$ and $1/\beta=B^{2}/8\pi p_{ext}\approx 1$ we represent 11-year variations of the magnetic flux tube in the photosphere for sunspots (see (28) or (27), in the form
@@
where $p_{ext}=6.53\cdot 10^{13}~erg/cm^{3}$ (see [49]), $T_{ext}=2.3\cdot 10^{6}~K$ and $\rho_{ext}=0.2~g/cm^{3}$ [49]; $c_{p}=2.5\Re=2.5p_{ext}/\rho_{ext}T_{ext}$ ($\Re$ is the gas constant in the equation of state $p_{ext}=\rho_{ext}\Re T_{ext}=nk_{B}T_{ext}$ [58]); $\mu=0.58$ is mean molecular weight of hydrogen in overshoot tachocline, $m_{p}\cong 1.67\cdot 10^{-24}~g$ is the proton mass: ”2018 CODATA Value: proton mass”, The NIST Reference on Constants, Units, and Uncertainty, 20 May 2019; $k_{B}=1.38\cdot 10^{-23}~J/K$; $(dQ/dt)_{1}\approx 29.7~erg\cdot cm^{-3}\cdot s^{-1}$ (see Eq. (18) in [64]) is the rate of radiative heating; $\alpha_{1}^{2}\approx 5.76$ [64, 67]; $\nabla_{e}\approx 0.4$ (see Table 2 in [61]); $H_{p}=\Re T_{ext}/g=p_{ext}/g\rho_{ext}=0.08R_{Sun}$ [58, 59, 60]; $a_{conv}^{min}\sim 3.7\cdot 10^{-4}H_{p}$ (see Eq. (24)).
@@
which are almost identical to the observational data of the known works (see about 1.4 km/s in [74, 75, 76, 77]), and at the minimum of the solar cycle
$(v_{rise})_{conv}^{min}\sim 0.14~km/s,$ (34)
which are almost identical to the observational data of the known works (see about $\sim 0.15~km/s$ in [78]), from which we obtain the “minimum” time of increase of convective heating $(dQ/dt)_{2}$ of the flux tube from the tachocline to the photosphere (see Eq. (15)), which corresponds to the equation
$(\tau_{d})_{conv}^{min}\approx 10^{6}~s\sim 10~days,$ (35)
where $p_{ext}=(nk_{B}T)_{min}=3.56\cdot 10^{13}~erg/cm^{3}$, $T_{ext}=10^{6}~K$, $\rho_{ext}\approx 0.25~g/cm^{3}$ (see Eq. (27)).
Ultimately, these Eq. (32) are the source of the first (fast in time) maximum of solar activity (e.g. sunspots) from the observed double maxima of 11-year cycles, while the second (two-year in time) maximum of sunspots is associated with observational data on the slope of Joy’s law (see Fig. 8b2,c on the right).
This raises an intriguing point regarding helioseismology. One of the first solutions in our section is the buoyancy of a virtually empty magnetic flux tube in strong magnetic fields of $\sim 10^{7}~G$ near the tachocline, determined by the thermomagnetic Ettingshausen-Nernst effect (see Apendix A in [14]), manifesting itself as “visible” sunspots on the solar surface. This is due to the fact that helioseismological inversions, which are well suited not only for studying the rotation and structure of the Sun, can sometimes be well suited due to the mediated effects of very strong magnetic fields at empty magnetic flux tubes. Let us repeat once again that it is the existence of empty magnetic flux tubes that allows photons of axion origin to “rush” past the tachocline to the photosphere both at high speed (see $\sim 1.4~km/s$ in [74, 75, 76, 77]) and at short times (see $\sim 1.3~days$ for Eqs. (31)-(32))! Thus, they can unequivocally confirm the presence of a magnetic field in the solar tachocline exceeding $10^{5}~G$ to, for example, $10^{7}~G$. On the other hand, it is the existence of “visible” sunspots in such strong fields that causes the dynamo effect to completely disappear.
The answer, surprisingly, is quite simple. It means that the “measured” velocity (maximum about $\sim 1.4~km/s$ [74, 75, 76, 77] and minimum about $\sim 0.15~km/s$ in [78]) of magnetic flux tubes, which are associated with structural changes, for example, in the form of cross-sections of these tubes at the maximum ($\sim 10^{22}~Mx$) or minimum ($5\cdot 10^{21}~Mx$) of the “measured” magnetic fluxes [71, 72], can be easily determined at the maximum $\sim 10^{7}~G$ or minimum $\sim 10^{6}~G$ values of the magnetic fields (see Eqs. (22)-(23)). The modernity of this considered helio-seismological inversion remains to this day!
In other words, the modernity of this well-thought-out helioseismological inversion remains to this day!
@@
Figure 6: Topological effects of secondary magnetic reconnection in the lower (left) or upper (right) layers of the magnetic tube. Here, the unipolar part of the $\Omega$-loop is rearranged at its base, compressing the $\Omega$-loop (blue lines) and, as a consequence, organizing a free O-loop (green lines) (a,e) to form (blue-yellow lines) the first part of the secondary magnetic reconnection (b,f). The yellow lines show the movement of matter leading to the loop connection “leg” (see analogous Fig. 4 in [52]). If the O-loop (primary magnetic reconnection, green lines) can randomly have different directions of magnetic fields, then the secondary magnetic reconnection can generate loop “legs” in different layers, for example, in lower layers (c and d as an analog of Fig. 4b) and upper layers (e and f as an analog of Fig. 4 in [79]).